OpenStudy (anonymous):

find c if a =5 and b=12

OpenStudy (anonymous):



OpenStudy (jdoe0001):

use the pythagorean theorem \(\bf c^2=a^2+b^2\implies c=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

so its 13

OpenStudy (jdoe0001):

so its 13 \(\large \checkmark\)